OpenStudy (anonymous):

Tell whether each function is linear. If so, graph the function. 5+3y=8

OpenStudy (katrinakaif):

How can we tell if an equation is linear or not?

OpenStudy (anonymous):

it can't be multiplied

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

think about what the powers must be

OpenStudy (anonymous):

it is linear but im not sure how to find the plots to graph it

OpenStudy (anonymous):

easiest way - is to put this in slope intercet form y = mx +b yint is 0, b - that is point #1 then from that point rise/run for your slope note - if neg slope down/run

OpenStudy (anonymous):

the other way is finding points by table -random values for x - plug in for values of y and then plot those points - connect and that is your graph i still prefer my first method.

OpenStudy (anonymous):

so put your equation in point slope form

OpenStudy (anonymous):

yes i do nderstand this is a horiztonal line, but the method still works. :)

OpenStudy (anonymous):

im not sure how to put it in point slope form

OpenStudy (anonymous):

reearrange what you were given to make it look like Y = MX + B

OpenStudy (anonymous):

so 3y=8+5

OpenStudy (anonymous):

SUBTRACT 5 FROM BOTH SIDES SO, 3Y = 8-5....

OpenStudy (anonymous):

3y=8+5 -5 -5 so 3y=3

OpenStudy (anonymous):

YES....NEXT

OpenStudy (anonymous):

yes so what should I do next to find the plots to graph

OpenStudy (anonymous):

there is NO coefficient in front of Y Y = MX + B

OpenStudy (anonymous):

SO YOU SHOULD BE AT Y = 1, DID YOU FOLLOW?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because I divided

OpenStudy (anonymous):

with that, there is no x term - so there is no slope..

OpenStudy (anonymous):

o ok thanks

OpenStudy (anonymous):

so all you have is a y- intecept that is 0,b or 0,1 in our case

OpenStudy (anonymous):